# Constructing an Optimal Behavior Basis for the Option Keyboard

### v1

@@ 4 Constructing an Optimal Behavior Basis
In this section, we introduce Option Keyboard Basis (OKB), a novel method to solve the goal introduced in the previous section. The OKB learns a set of base policies, $\Pi_{k}$, and a meta-policy, $\omega$, such that the induced OK policy $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is provably optimal w.r.t. any given task $\bm{\mathbf{w}}\in\mathcal{W}$. OKB’s pseudocode is shown in Alg. 1.
The algorithm starts with a single base policy—optimized to solve an arbitrary initial task; e.g., $\bm{\mathbf{w}}{=}[1/d,...,1/d]^{\top}$—in its set of policies $\Pi$ (lines 2–3). OKB’s initial partial CCS, $\Psi^{\text{OK}}$, and weight support set, $\mathcal{W}^{\text{sup}}$ are initialized as empty sets (lines 4–5). $\mathcal{W}^{\text{sup}}$ will store the weights of tasks the meta-policy has been trained on. At each iteration $k$, OKB carefully selects the weight vectors on which its meta-policy, $\omega$, is trained so that the policies expressible by $\Psi^{\text{OK}}$ iteratively approximate a CCS. This process is implemented by OK-LS (Alg. 2), an algorithm inspired by the SFs Optimistic Linear Support (SFOLS) method (Alegre et al., , 2022) but that operates over a meta-policy $\omega$ rather than the space of base policies.
@@
1:  Input: MPC $(\mathcal{S},\mathcal{A},p,\gamma,\mu)$ with features $\bm{\mathbf{\phi}}(s,a){\in}\mathbb{R}^{d}$.
[…]
3:  $\Pi_{0}\leftarrow\{\pi_{\bm{\mathbf{w}}}\};\Psi^{\text{base}}_{0}\leftarrow\{\bm{\mathbf{\psi}}^{\pi_{\bm{\mathbf{w}}}}\}$
4:  $\Psi^{\text{OK}}_{0}\leftarrow\{\};\mathcal{W}^{\text{sup}}_{0}\leftarrow\{\}$
5:  Initialize meta-policy $\omega_{0}$
6:  for $k=0,...,\infty$ do
7:     $\triangleright$ Update meta-policy $\omega$
8:     $\omega_{k+1},\Psi^{\text{OK}}_{k+1},\mathcal{W}^{\text{sup}}_{k+1}\leftarrow\texttt{OK-LS}(\omega_{k},\Psi^{\text{OK}}_{k},\mathcal{W}^{\text{sup}}_{k},\Pi_{k})$
9:     $\mathcal{C}\leftarrow\{\}$
10:     for $\bm{\mathbf{w}}\in\texttt{{CornerW}}(\Psi^{\text{base}}_{k})\cup\texttt{{CornerW}}(\Psi^{\text{OK}}_{k+1})$ do
11:        $\triangleright$ Check if task $\bm{\mathbf{w}}$ is solvable with OK and $\Pi_{k}$
12:        if $\pi^{*}_{\bm{\mathbf{w}}}$ is not in $\Pi^{\textup{OK}}(\Pi_{k})$ then
13:           $\mathcal{C}\leftarrow\mathcal{C}\cup\{\bm{\mathbf{w}}\}$
14:        end if
15:     end for
16:     if $\mathcal{C}$ is empty then
17:        $\triangleright$ Found optimal basis $\Pi$ and meta-policy $\omega$
18:        Return $\omega_{k+1},\Pi_{k},\Psi^{\text{OK}}_{k+1}$
19:     else
20:        $\triangleright$ Learn a new base policy
21:        $\bm{\mathbf{w}}\leftarrow$ select a task from $\mathcal{C}$
22:        $\pi_{\bm{\mathbf{w}}},\bm{\mathbf{\psi}}^{\pi_{\bm{\mathbf{w}}}}\leftarrow\texttt{NewPolicy}(\bm{\mathbf{w}})$
23:        $\Pi_{k+1},\leftarrow\Pi_{k}\cup\{\pi_{\bm{\mathbf{w}}}\};\Psi^{\text{base}}_{k+1}\leftarrow\Psi^{\text{base}}_{k}\cup\{\bm{\mathbf{\psi}}^{\pi_{\bm{\mathbf{w}}}}\}$
24:        $\Pi_{k+1},\Psi^{\text{base}}_{k+1}{\leftarrow}\texttt{RemoveDominated}(\Pi_{k+1},\Psi^{\text{base}}_{k+1})$
25:     end if
26:  end for
@@
1:  Input: Meta-policy $\omega$, partial CCS $\Psi^{\text{OK}}$, weight support $\mathcal{W}^{\text{sup}}$, base policies $\Pi_{k}$.
2:  while True do
3:     $\mathcal{W}^{\text{corner}}\leftarrow\texttt{CornerW}(\Psi^{\text{OK}})\setminus\mathcal{W}^{\text{sup}}$
4:     if $\mathcal{W}^{\text{cw}}$ is empty then
5:        Return $\omega,\Psi^{\text{OK}},\mathcal{W}^{\text{sup}}$
6:     end if
7:     $\mathcal{W}^{\text{sup}}\leftarrow\mathcal{W}^{\text{sup}}\cup\mathcal{W}^{\text{corner}}$
8:     $\omega\leftarrow\texttt{TrainOK}(\omega,\mathcal{W}^{\text{sup}},\Pi_{k})$triangleright (Alg. 3)
9:     $\Psi^{\text{OK}}\leftarrow\{\bm{\mathbf{\psi}}^{\pi^{\text{OK}}_{\omega}(\cdot,\bm{\mathbf{w}};\Pi_{k})}\mid\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}\}$
10:     $\Psi^{\text{OK}}\leftarrow\texttt{RemoveDominated}(\Psi^{\text{OK}})$
11:  end while
@@
Next, in lines 10–15, OKB identifies a set of candidate tasks, $\mathcal{C}$, that the OK cannot optimally solve with its current set of policies, $\Pi_{k}$—i.e., tasks for which $\pi_{\bm{\mathbf{w}}}^{*}$ is not included in $\Pi^{\text{OK}}(\Pi_{k})$. We discuss how to check this condition in Section 4.1 If $\mathcal{C}$ is empty (line 16), the OKB terminates and returns a meta-policy ($\omega$) and base policies ($\Pi_{k}$) capable of ensuring that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$. This is due to Thm. 4.3, which we discussed later. If $\mathcal{C}$ is not empty, OKB selects a task from it and adds a new corresponding optimal base policy to its behavior basis, $\Pi_{k}$ (lines 20–23). In line 24, RemoveDominated removes redundant policies—policies that are not strictly required to solve at least one task.
@@
In Algorithms 1 and 2, the function $\texttt{CornerW}(\Psi)$ takes a set of SF vectors, $\Psi$, as input and returns a set of corner weights (see App. E.1 for its mathematical definition). Intuitively, both Alg. 1 and Alg. 2 rely on the fact that (i) a task $\bm{\mathbf{w}}\in\mathcal{W}$ that maximizes $\Delta(\bm{\mathbf{w}})\triangleq v^{*}_{\bm{\mathbf{w}}}-\max_{\pi\in\Pi_{k}}v^{\pi}_{\bm{\mathbf{w}}}$ is guaranteed to be a corner weight, and (ii) if OKB has an optimal policy for all corner weights, then it must have identified a CCS. These results were demonstrated in prior work on constructing a CCS (Roijers et al., , 2013; Alegre et al., , 2022) and can also be proven using the fundamental theorem of linear programming.
@@
OK-LS (Alg. 2) trains the meta-policy $\omega$, on selected tasks $\mathcal{W}^{\text{cw}}$ (line 3) using the base policies $\Pi_{k}$, so that is partial CCS, $\Psi^{\text{OK}}$, iteratively approximates a CCS. It stores the tasks $\bm{\mathbf{w}}$ it has already trained on, as well as the corner weights of the current iteration, in the weight support set, $\mathcal{W}^{\text{sup}}$ (line 7). If $\mathcal{W}^{\text{cw}}$ is empty in a given iteration, no tasks remain to be solved, and the algorithm returns the updated meta-policy. Otherwise, OK-LS adds the corner weights $\mathcal{W}^{\text{cw}}$ to $\mathcal{W}^{\text{sup}}$ and trains the meta-policy $\omega$ on the tasks $\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}$ using the TrainOK subroutine (Alg. 3). Finally, in line 9, OK-LS computes the SF vectors of the policies induced by OK for each $\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}$, updating its partial CCS, $\Psi^{\text{OK}}$. [^3] Note that to reduce the algorithm’s computational cost, SF vectors can be trained and computed only for the weight vectors in $\mathcal{W}^{\text{sup}}$ that are new in the current iteration (lines 8–9). In App. C, we discuss how we train $\omega$ using an actor-critic RL method (Alg. 3). Finally, note that to accelerate the learning of $\omega$, the corner weights in $\mathcal{W}^{\text{cw}}$ can be prioritized, similar to Alegre et al., (2022).
@@ 4.1 Condition for OK Optimality
In this section, we address the following question: Given an arbitrary task with reward function $r$ and a set of base policies $\Pi_{k}$, is $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$? In other words, does a meta-policy $\omega$ exist such that the OK policy, $\pi^{\text{OK}}{\omega}(\cdot;\Pi_{k})$, can represent the optimal policy $\pi_{r}^{*}$? Determining whether this holds is essential for implementing the OKB step in line 12.
@@
This proposition provides a sufficient condition for the existence of an optimal OK meta-policy $\omega$ for a given reward function $r$. Intuitively, the OK policy should be able to express all optimal actions through $\omega$ and $\Psi$. Next, we show how to verify this condition without requiring access to $q^{*}_{r}$.
@@
Let $\Pi_{k}$ be a set of base policies and $\omega^{*}$ be a meta-policy trained to convergence to solve a given task $r$. If $A^{\omega^{*}}{r}(s,a)>0$ for some $(s,a)$, then action $a$ is not expressible by $\pi^{\text{OK}}{\omega^{*}}(\cdot;\Pi_{k})$. Consequently, the OK with base policies $\Pi_{k}$ cannot represent an optimal policy for $r$.
For a detailed discussion on applying Thm. 4.2 in practical implementations of the OKB to verify whether $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$, see App. D.
@@ 4.2 Theoretical Results
OKB returns a set of base policies $\Pi_{k}$ and a meta-policy $\omega$ such that $|\Pi_{k}|\leq|\mathrm{CCS}|$ and $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is optimal for any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. This implies that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$; i.e., the OK is at least as expressive as a CCS.
This theorem states the main result of this paper: the proposed method, OKB (Alg. 1), achieves the primary goal introduced in Section 3. That is, OKB identifies a behavior basis, $\Pi_{k}$, that can be used by an option keyboard to optimally solve any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. Furthermore, since OKB allows optimal policies to be expressed with fewer base policies than a CCS, the total computational cost of identifying optimal solutions for all $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ is significantly reduced compared to GPI-based algorithms. This is because computing all policies needed to form a CCS becomes intractable as the number of reward features $d$ grows. As a result, the computational cost gains provided by OKB become increasingly pronounced as task complexity increases.
To state our next theoretical result, we first recall Thm. 2 by Barreto et al., (2017):
Theorem 4.4 (Barreto et al., (2017)).
Let $\Pi=\{\pi_{i}\}_{i=1}^{n}$ be a set of optimal policies w.r.t. tasks $\{\bm{\mathbf{w}}_{i}\}_{i=1}^{n}$ and let $\{\hat{\bm{\mathbf{\psi}}}^{\pi_{i}}\}_{i=1}^{n}$ be approximations to their SFs. Let $\bm{\mathbf{w}}\in\mathcal{W}$ be a task and let $|q^{\pi_{i}}_{\bm{\mathbf{w}}}(s,a)-\hat{q}^{\pi_{i}}_{\bm{\mathbf{w}}}(s,a)|\leq\epsilon\text{ for all }(s,a)\in\mathcal{S}\times\mathcal{A},\text{ and }\,\pi_{i}\in\Pi$. Let $\bm{\mathbf{\phi}}_{\mathrm{max}}\triangleq\max_{s,a}||\bm{\mathbf{\phi}}(s,a)||$. Then, it holds that, for all $(s,a)\in\mathcal{S}\times\mathcal{A}$:
$q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{GPI}}}_{\bm{\mathbf{w}}}(s,a)\leq\frac{2}{1-\gamma}\left(\bm{\mathbf{\phi}}_{\mathrm{max}}\min_{i}||\bm{\mathbf{w}}-\bm{\mathbf{w}}_{i}||+\epsilon\right).$
This theorem describes the optimality gap of GPI policies, how it depends on the available base policies, and how it is affected by function approximation errors. It does not, however, characterize the optimality gap of option keyboards, which generalize GPI policies. We introduce and highlight two generalizations of this theorem: (1) Since the OK generalizes GPI policies, it follows that $q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{OK}}_{\omega}}_{\bm{\mathbf{w}}}(s,a)\leq q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{GPI}}}_{\bm{\mathbf{w}}}(s,a)$; and (2) when Eq. (10) (Prop. 4.1) is satisfied, $q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{OK}}_{\omega}}_{\bm{\mathbf{w}}}(s,a)=0$. That is, the OK can completely avoid optimality gaps due to approximation errors in the base policies’ SFs. Hence, not only are the OK and OKB more expressive than GPI (Thm. 4.3), but they also naturally lead to transfer learning strategies that are significantly more robust to approximation errors than GPI.
Theorem 4.5.
@@
This theorem precisely characterizes a class of non-linear tasks that can be optimally solved by an OK using the behavior basis learned by OKB. Intuitively, OKB can solve any task whose optimal policy can be constructed by alternating between the optimal policies for tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$, even when the task itself is non-linear. Importantly, existing methods that compute a CCS can only handle linearly expressible tasks, whereas OKB extends this capability to a broader class of problems. Furthermore, OKB achieves this without having to learn all policies in the CCS (Thm. 4.3). We further discuss the properties of OK under non-linear reward functions in App. B.

### v2

@@ 4 Constructing an Optimal Behavior Basis
We now introduce Option Keyboard Basis (OKB), a novel method to solve the goal introduced in the previous section. The OKB (Alg. 1) learns a set of base policies, $\Pi_{k}$, and a meta-policy, $\omega$, such that the induced OK policy $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is provably optimal w.r.t. any given task $\bm{\mathbf{w}}\in\mathcal{W}$.
The algorithm starts with a single base policy optimized to solve an arbitrary initial task (e.g., $\bm{\mathbf{w}}{=}[1/d,...,1/d]^{\top}$) in its set of policies $\Pi$ (lines 1–2). OKB’s initial partial CCS, $\Psi^{\text{OK}}$, and weight support set, $\mathcal{W}^{\text{sup}}$, are initialized as empty sets (line 3). $\mathcal{W}^{\text{sup}}$ will store the weights of tasks the meta-policy has been trained on. At each iteration $k$, OKB carefully selects the weight vectors on which its meta-policy, $\omega$, will be trained so that the policies expressible by $\Psi^{\text{OK}}$ iteratively approximate a CCS. This process is implemented by OK-LS (Alg. 2), an algorithm inspired by the SFs Optimistic Linear Support (SFOLS) method (Alegre+2022) but that operates over a meta-policy $\omega$ rather than the space of base policies.
1
Input: MPC with features $\bm{\mathbf{\phi}}(s,a)\in\mathbb{R}^{d}$
2
[…]
4 $\Pi_{0}\leftarrow\{\pi_{\bm{\mathbf{w}}}\};\Psi^{\text{base}}_{0}\leftarrow\{\bm{\mathbf{\psi}}^{\pi_{\bm{\mathbf{w}}}}\}$
5 $\Psi^{\text{OK}}_{0}\leftarrow\{\};\mathcal{W}^{\text{sup}}_{0}\leftarrow\{\}$
6 Initialize meta-policy $\omega_{0}$
7
8for $k=0,1,2,\dots$ do
  $\triangleright$ Update meta-policy $\omega$
9  $\omega_{k+1},\Psi^{\text{OK}}_{k+1},\mathcal{W}^{\text{sup}}_{k+1}\leftarrow\texttt{OK-LS}(\omega_{k},\Psi^{\text{OK}}_{k},\mathcal{W}^{\text{sup}}_{k},\Pi_{k})$
10
11 $\mathcal{C}\leftarrow\{\}$
12
13 for $\bm{\mathbf{w}}\in\textnormal{{CornerW}}(\Psi^{\text{base}}_{k})\cup\textnormal{{CornerW}}(\Psi^{\text{OK}}_{k+1})$ do
     $\triangleright$ Check if task $\bm{\mathbf{w}}$ is solvable with OK and $\Pi_{k}$
14     if $\pi^{*}_{\bm{\mathbf{w}}}$ is not in $\Pi^{\textup{OK}}(\Pi_{k})$ then
15        $\mathcal{C}\leftarrow\mathcal{C}\cup\{\bm{\mathbf{w}}\}$
16
17 if $\mathcal{C}$ is empty then
     $\triangleright$ Found optimal basis $\Pi$ and meta-policy $\omega$
18     return $\omega_{k+1},\Pi_{k},\Psi^{\text{OK}}_{k+1}$
19 else
     $\triangleright$ Learn a new base policy
20     $\bm{\mathbf{w}}\leftarrow$ select a task from $\mathcal{C}$
21     $\pi_{\bm{\mathbf{w}}},\bm{\mathbf{\psi}}^{\pi_{\bm{\mathbf{w}}}}\leftarrow\texttt{NewPolicy}(\bm{\mathbf{w}})$
22     $\Pi_{k+1}\leftarrow\Pi_{k}\cup\{\pi_{\bm{\mathbf{w}}}\}$
23     $\Psi^{\text{base}}_{k+1}\leftarrow\Psi^{\text{base}}_{k}\cup\{\bm{\mathbf{\psi}}^{\pi_{\bm{\mathbf{w}}}}\}$
24     $\Pi_{k+1},\Psi^{\text{base}}_{k+1}\leftarrow\texttt{RemoveDominated}(\Pi_{k+1},\Psi^{\text{base}}_{k+1})$
25
@@ 22     $\Pi_{k+1}\leftarrow\Pi_{k}\cup\{\pi_{\bm{\mathbf{w}}}\}$
1
2
Input: Meta-policy $\omega$, partial CCS $\Psi^{\text{OK}}$, weight support $\mathcal{W}^{\text{sup}}$, base policies $\Pi_{k}$.
3
4while True do
5  $\mathcal{W}^{\text{corner}}\leftarrow\texttt{CornerW}(\Psi^{\text{OK}})\setminus\mathcal{W}^{\text{sup}}$
6  if $\mathcal{W}^{\text{cw}}$ is empty then
7     return $\omega,\Psi^{\text{OK}},\mathcal{W}^{\text{sup}}$
8 $\mathcal{W}^{\text{sup}}\leftarrow\mathcal{W}^{\text{sup}}\cup\mathcal{W}^{\text{corner}}$
9
 $\omega\leftarrow\texttt{TrainOK}(\omega,\mathcal{W}^{\text{sup}},\Pi_{k})$ $\triangleright$ (Alg. 3)
10
11 $\Psi^{\text{OK}}\leftarrow\{\bm{\mathbf{\psi}}^{\pi^{\text{OK}}_{\omega}(\cdot,\bm{\mathbf{w}};\Pi_{k})}\mid\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}\}$
12
13 $\Psi^{\text{OK}}\leftarrow\texttt{RemoveDominated}(\Psi^{\text{OK}})$
14
@@ 13 $\Psi^{\text{OK}}\leftarrow\texttt{RemoveDominated}(\Psi^{\text{OK}})$
Next, in lines 8–10, OKB identifies a set of candidate tasks, $\mathcal{C}$, that the OK cannot optimally solve with its current set of policies, $\Pi_{k}$; i.e., a set of tasks for which $\pi_{\bm{\mathbf{w}}}^{*}$ is not included in $\Pi^{\text{OK}}(\Pi_{k})$. We discuss how to check this condition in Section 4.1 If $\mathcal{C}$ is empty (line 11), the OKB terminates and returns a meta-policy ($\omega$) and base policies ($\Pi_{k}$) capable of ensuring that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$. This is due to Thm. 4.3, which we discuss later. If $\mathcal{C}$ is not empty, OKB selects a task from it and adds a new corresponding optimal base policy to its behavior basis, $\Pi_{k}$ (lines 14–17). In line 18, RemoveDominated removes redundant policies—policies that are not strictly required to solve at least one task.
@@
In Algorithms 1 and 2, the function $\texttt{CornerW}(\Psi)$ takes a set of SF vectors, $\Psi$, as input and returns a set of corner weights (see App. F.1). Intuitively, both Alg. 1 and Alg. 2 rely on the fact that (i) a task $\bm{\mathbf{w}}\in\mathcal{W}$ that maximizes $\Delta(\bm{\mathbf{w}})\triangleq v^{*}_{\bm{\mathbf{w}}}-\max_{\pi\in\Pi_{k}}v^{\pi}_{\bm{\mathbf{w}}}$ is guaranteed to be a corner weight, and (ii) if OKB has an optimal policy for all corner weights, then it must have identified a CCS. These results were shown in prior work on constructing a CCS (Roijers2016; Alegre+2022) and can also be proven using the fundamental theorem of linear programming.
@@
OK-LS (Alg. 2) trains the meta-policy, $\omega$, on selected tasks $\mathcal{W}^{\text{cw}}$ (line 2) using the base policies $\Pi_{k}$, so that is partial CCS, $\Psi^{\text{OK}}$, iteratively approximates a CCS. It stores the tasks $\bm{\mathbf{w}}$ it has already trained on, as well as the corner weights of the current iteration, in the weight support set, $\mathcal{W}^{\text{sup}}$ (line 5). If $\mathcal{W}^{\text{cw}}$ is empty in a given iteration, no tasks remain to be solved, and the algorithm returns the updated meta-policy. Otherwise, OK-LS adds the corner weights $\mathcal{W}^{\text{cw}}$ to $\mathcal{W}^{\text{sup}}$ and trains the meta-policy $\omega$ on the tasks $\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}$ using the TrainOK subroutine (Alg. 3). Finally, in line 7, OK-LS computes the SF vectors of the policies induced by OK for each $\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}$, updating its partial CCS, $\Psi^{\text{OK}}$. [^3] Notice that SFs only need to be computed for vectors newly added to $\mathcal{W}^{\text{sup}}$ in the current iteration (lines 6-7). In App. D, we discuss how to train $\omega$ using an actor-critic RL method (Alg. 3). Finally, note that to accelerate the learning of $\omega$, the corner weights in $\mathcal{W}^{\text{cw}}$ can be prioritized, similar to Alegre+2022.
@@ 4.1 Condition for OK Optimality
In this section, we address the following question: Given an arbitrary task with reward function $r$ and a set of base policies $\Pi_{k}$, is $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$? In other words, does a meta-policy $\omega$ exist such that the OK policy, $\pi^{\text{OK}}{\omega}(\cdot;\Pi_{k})$, can represent the optimal policy $\pi_{r}^{*}$? Determining whether this holds is essential for implementing the OKB step in line 9.
@@
This proposition provides a sufficient condition for the existence of an optimal OK meta-policy $\omega$ for a given reward function $r$. Intuitively, it implies that the OK policy should be able to express all optimal actions through $\omega$ and $\Psi$. Next, we show how to verify this condition without requiring access to $q^{*}_{r}$.
@@
Let $\Pi_{k}$ be a set of base policies and $\omega^{*}$ be a meta-policy trained to convergence to solve a given task $r$. If $A^{\omega^{*}}_{r}(s,a)>0$ for some $(s,a)$, then action $a$ is not expressible by $\pi^{\text{OK}}_{\omega^{*}}(\cdot;\Pi_{k})$. Consequently, the OK with base policies $\Pi_{k}$ cannot represent an optimal policy for $r$.
Note that the procedure for verifying whether the OK is optimal for a given task $r$ (Thm. 4.2) is a sufficient, but not necessary, condition to guarantee optimality. It serves two key purposes: (i) to avoid training on tasks $\omega$ whose optimal policies have either already been learned or that can be reconstructed from $\Pi_{k}$ and $\omega$, and (ii) to prioritize which tasks to train on in order to accelerate learning. In other words, the procedure for assessing OK optimality is primarily a tool to reduce training effort; our method does not rely on an exact test to ensure the correctness of the behavior basis it constructs. For a detailed discussion of how Thm. 4.2 can be applied in practical implementations of OKB to verify whether $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$, see App. E.
@@ 4.2 Theoretical Results
OKB returns a set of base policies, $\Pi_{k}$, and a meta-policy, $\omega$, such that $|\Pi_{k}|\leq|\mathrm{CCS}|$ and $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is optimal for any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. This implies that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$; i.e., the OK matches or exceeds the expressiveness of a CCS and ensures zero-shot optimality for all $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$.
This is a key result of this paper: OKB can provably identify a behavior basis, $\Pi_{k}$, that enables an option keyboard to optimally solve any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. Furthermore, it is capable of ensuring zero-shot optimality using a set of base policies potentially smaller than a CCS. This is the first method to offer such a guarantee—until now, all existing approaches required the number of base policies to be strictly equal to the size of the CCS. This limitation posed a major scalability challenge: computing a full CCS becomes intractable as the number of reward features $d$ grows, since its size can grow exponentially with dd—counting corner weights is equivalent to Vertex Enumeration (Roijers2016). OKB overcomes this bottleneck by avoiding the need to construct a full CCS, thereby reducing the total cost of computing optimal policies for all $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. See below for more details.
Proposition 4.4.
Let $\Pi_{k}$ be the set of policies learned by OKB (Alg. 1) when solving tasks that are linearly expressible in terms of reward features $\bm{\mathbf{\phi}}(s,a)\in\mathbb{R}^{d}$. Let $r$ be an arbitrary reward function that is non-linear with respect to $\bm{\mathbf{\phi}}$. Suppose the optimal policy for $r$ can be expressed by alternating, as a function of the state, between policies in the CCS induced by $\bm{\mathbf{\phi}}$. That is, suppose that for all $s\in\mathcal{S}$, there exists a policy $\pi^{*}_{\bm{\mathbf{w}}}$ (optimal for some $\bm{\mathbf{w}}\in\mathcal{W}$) such that $\pi^{*}_{r}(s)=\pi^{*}_{\bm{\mathbf{w}}}(s)$. Then, the OK can represent the optimal policy for $r$ using the set of base policies $\Pi_{k}$; i.e., $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$.
This deepens the theoretical understanding of the OKB by showing it is strictly more expressive than existing methods. In particular, it allows us to characterize (1) conditions under which the OKB matches the expressiveness of a CCS while relying on strictly fewer policies; and (2) a broad class of non-linear tasks that can be optimally solved by an OK using the behavior basis learned by OKB.
First, Prop. 4.4 establishes the optimality of the OKB in transfer learning scenarios where related tasks can be solved by reusing parts of existing solutions—that is, the widely studied setting in which transfer is possible due to shared structure across tasks. Concretely, the OKB is guaranteed to optimally solve, in a zero-shot manner, tasks whose solutions decompose into sequences of sub-policies. Building on this result and Thm. 4.3, we formally characterize, in App. A.5, the conditions under which the OKB is as expressive as a CCS while relying on strictly fewer policies.
@@
Second, Proposition 4.4 precisely characterizes a class of non-linear tasks that can be optimally solved by an OK using the behavior basis learned by OKB. Intuitively, OKB can solve any task whose optimal policy can be constructed by alternating between the optimal policies for tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$, even when the task itself is non-linear. This implies, importantly, that while existing methods that compute a CCS are limited to linearly expressible tasks, OKB can handle a broader class of non-linear tasks. Furthermore, as discussed above, the OKB achieves this without having to learn all policies in the CCS. We further discuss the properties of OK under non-linear reward functions in App. C.
To state our next theoretical result, we first recall Thm. 2 by Barreto+2017:
Theorem 4.5 (Barreto+2017).
[…]
Thm 4.5 describes the optimality gap of GPI policies, how it depends on the available base policies, and how it is affected by function approximation errors. It does not, however, characterize the optimality gap of option keyboards, which generalize GPI policies. We introduce and highlight two generalizations of this theorem: (1) Since the OK generalizes GPI policies, it follows that $q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{OK}}_{\omega}}_{\bm{\mathbf{w}}}(s,a)\leq q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{GPI}}}_{\bm{\mathbf{w}}}(s,a)$; and (2) when Eq. (8) (Prop. 4.1) is satisfied, $q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{OK}}_{\omega}}_{\bm{\mathbf{w}}}(s,a)=0$. That is, OK can entirely avoid optimality gaps from approximation errors in the base policies’ SFs. Thus, OK and OKB are not only more expressive than GPI (Thm. 4.3) but also naturally lead to transfer learning strategies that are significantly more robust to approximation errors than GPI.